Solve the differential equation dy/dx +[x/1-x^2]y= x√y. y(0) =1
1
Expert's answer
2020-05-26T20:11:28-0400

the given equation can be written as:

"\frac{y'}{\sqrt{y}}+ \frac{x\sqrt{y}}{(1-x^2)} = x"


let "v=2\sqrt{y}"

then, "v'=\frac{y'}{\sqrt{y}}"

substituting above, we get,

"v'+ \frac{xv}{2(1-x^2)} = x" ........(1)


comparing with v'+Pv=Q where P and Q are functions of x

P= "\frac{x}{1-x^2}" and Q= x


integrating factor= "e^{\int{Pdx}}= e^{\int{\frac{x}{(1-x^2)}dx}} = (1-x^2)^{-1/4}"


multiplying (1) by integrating factor,

"v. (1-x^2)^{-1/4} = \int{(1-x^2)^{-1/4}.x}dx"


"v. (1-x^2)^{-1/4} = -\frac{2}{3}(1-x^2)^{3/4} + c"


substituting value of v,


"2\sqrt{y}. (1-x^2)^{-1/4} = -\frac{2}{3}(1-x^2)^{3/4} + c"


"2\sqrt{y}= -\frac{2}{3}(1-x^2) + c(1-x^2)^{1/4}"


now, y(0)=1


"2=\frac{-2}{3}+c \\

c=\frac{8}{3}"


"2\sqrt{y}= -\frac{2}{3}(1-x^2) + \frac{8}{3}(1-x^2)^{1/4}"